OpenStudy (anonymous):

will fan and give medal!!! determine whether y varies directly with x. if so, find the constant of variation k. 3y=-7x-18

OpenStudy (gorv):

for direct variation equation should of tjis type y=cx where c is constant of vavriation

OpenStudy (anonymous):

i dont understand at all D: I'm sorry. :/

OpenStudy (gorv):

like if y vary than value of x should vary in a proportion

OpenStudy (anonymous):

oooh, alright. so pretty much much 1/3=2/6 type of thing?

OpenStudy (anonymous):

so would i have to solve the equation twice. 1 for y and 1 for x?

OpenStudy (gorv):

only difference is it will nultiplied with wome contant like y=2x x=1 than y=2 x=2 than y= 4 x=9 than y=18 hope you got it

OpenStudy (anonymous):

alright, can you explain how to solve the equation part? I'm sorry.

OpenStudy (gorv):

no no like x have value 1 than put value of x and get value of y

OpenStudy (gorv):

3y=-7x-18 like x= 0 put it in equation 3y=-7*0-18 3y=-18 y=-6

OpenStudy (gorv):

but in it you have to -18 so it is not directly proportional

OpenStudy (anonymous):

so the x would have to equal the -18? or? haha, I'm sorry, I'm really confused

OpenStudy (gorv):

loll....here we can put any value for x WHATEVER YOU WANT but after putting value of x -7*x -18 you have to subtract 18 from the value obtained

OpenStudy (gorv):

llet x=2 y=-7*2-18 y=-14-18=-32 for directly proportionality nothing should be added or subtracted from x term but here 18 is subtracted so it is not directly vary

OpenStudy (anonymous):

alright, so what your saying is when another number is subtracted from it, it does not directly vary?

OpenStudy (gorv):

yeah subtracted or added

OpenStudy (gorv):

y=2x y=100x y=x/478748 y=x/677 all vary directly

OpenStudy (gorv):

but y=2x+3 y=100x-20 not vary directly

OpenStudy (anonymous):

oooh, alright! that makes more sense haha. so the answer would be no?

OpenStudy (gorv):

yepppppppppppppp

OpenStudy (anonymous):

thank you so much!

OpenStudy (gorv):

welcm lol :)